The sum of two numbers is 80. The larger number is 8 less than 3 times the smaller number. find the smaller number
Answer by malakumar_kos@yahoo.com(315)


The sum of two numbers is 80. The larger number is 8 less than 3 times the smaller number. find the smaller number

Let the smaller number be x
the larger number = 3x- 8
sum of the two no's = 80
x +(3x-8) = 80
x+3x = 80+8
4x = 88 or x = 88/4 = 22
the smaller no = 22 and the larger no = 58